OpenStudy (anonymous):

Solve. Express your answer in simplest form. 4 3/7÷6/7 A. 13/49 B. 6/31 C. 5 1/6 D. 4 2/7

OpenStudy (anonymous):

B

OpenStudy (igreen):

@TheOnlyDarkKnight Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her participation. Please read the \(\bf\color{blue}{\underline{{\href{http:///openstudy.com/code-of-conduct}{Code-of-Conduct}}}}\) to learn more.

OpenStudy (igreen):

Convert 4 3/7 to an improper fraction. 7 * 4 + 3 = ?

OpenStudy (anonymous):

If we are using pemdas then it will be it will be 4 3/7 7x3= 21+4= 25/7

OpenStudy (igreen):

We multiply the denominator to the whole number and add it to the numerator.

OpenStudy (igreen):

\(\dfrac{7 \times 4 + 3}{7}\)

OpenStudy (anonymous):

31/7

OpenStudy (igreen):

Correct, now dividing by 6/7 is the same as multiplying by its reciprocal, 7/6. So multiply: \(\dfrac{31}{7} \times \dfrac{7}{6}\)

OpenStudy (anonymous):

@iGreen it says "we do not encourage giving answers"

OpenStudy (anonymous):

217/42

OpenStudy (igreen):

Correct, and that simplifies to: 31/6 Can you convert that to mixed number form?

OpenStudy (igreen):

@TheOnlyDarkKnight It says "Give Help, NOT ANSWERS"

OpenStudy (igreen):

Plus, all the moderators of the site have agreed that if there is no participation of the asker, you are breaking the Code-of-Conduct.

OpenStudy (anonymous):

so its c 5 1/6

OpenStudy (igreen):

Correct

OpenStudy (anonymous):

"will >>>>>>encourage<<<<< and guide those needing help, and not just give them an answer"

OpenStudy (igreen):

\(\bbox [10pt, lime, border:5pt solid black]{\huge\cal\color{red}\succ\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\prec}}}\) You can give medals by clicking 'Best Response'. @Allsia1245